OpenStudy (anonymous):

Parallel and Perpendicular Lines: 1. Write the equation of the line which passes though (-4, 2) and is parallel to y = -x + 6 in slop-intercept form. 2. Write the equation of the line which passes though (2, -3) and is perpendicular to y = 4x + 7 in slope-intercept form.

OpenStudy (texaschic101):

1. If you are looking for a parallel line, the slope will be the same. So, in this case, the slope is -1. y - y1 = m(x - x1) using slope(m) - 1 and points (-4,2) y - 2 = -1(x - (-4) y - 2 = -1(x + 4) y - 2 = -x - 4 y = -x - 4 + 2 y = -x - 2 <-- this is it :) 2. If you are looking for perpendicular, the slope will be the negative reciprocal. So, in this case, the slope is -1/4. y - y1 = m(x - x1) using slope -1/4 and points (2,-3) y - (-3) = -1/4(x - 2) y + 3 = -1/4x + 2/4 y = -1/4x + 2/4 - 3 y = -1/4x + 2/4 - 12/4 y = -1/4x - 10/4 y = -1/4x - 5/2 <-- this is it :)

OpenStudy (anonymous):

eq. of line parallel to y=-x+6 is y=-x+c because it passes through (-4,2) 2=-(-4)+c 2=4+c c=2-4=-2 reqd. eq is y=-x-2 or x+y+2=0

OpenStudy (anonymous):

slope form is y=-x-2